OpenStudy (anonymous):

Write an inequality to represent the problem. Then solve the inequality by writing the pairs which solve it. Find all sets of two consecutive positive odd integers whose sum is less than or equal to 18.

OpenStudy (anonymous):

10

OpenStudy (anonymous):

\(0<n+n+2\leq 18\) might work

OpenStudy (anonymous):

it is easy enough to solve smallest odd integer is 1, so 1,3 works 3,5 5,7 etc

OpenStudy (anonymous):

well there are no too many more, might as well complete it

OpenStudy (anonymous):

Mmm it could, I just came up with 2n + 1 <= 18. Getting a second opinion coz a lot of the things I've tried have a lot of parings, don't want one with many :P

OpenStudy (anonymous):

they have to be positive, which is why i put the zero on the left

OpenStudy (anonymous):

also if they are consecutive ODD integers, then if one is \(n\) the other is \(n+2\)

OpenStudy (anonymous):

7,9 is the last one, because \(9+11=20\) is too large

OpenStudy (anonymous):

Ooohhh, okay! I get it now, thank you so much (: